OpenStudy (albert0898):

Find d2y/dx2 in terms of x and y. x^2y^2 - 16x = 4

OpenStudy (anonymous):

did you find \(y'\) first?

OpenStudy (anonymous):

by 'implicit differentiation"?

OpenStudy (albert0898):

y' would be \[\frac{2\left(2x+1\right)\sqrt{\left(x\right)^2}}{x^3\sqrt{4x+1}}\] ?

OpenStudy (anonymous):

i see no square roots in your question

OpenStudy (anonymous):

you posted this :\[x^2y^2 - 16x = 4\]

OpenStudy (albert0898):

What I did was put everything in terms of y? Would that be wrong?

OpenStudy (anonymous):

yes it would

OpenStudy (albert0898):

So, how would I approach this problem?

OpenStudy (anonymous):

you familiar with implicit differentiation?

OpenStudy (albert0898):

Yes

OpenStudy (anonymous):

use that

OpenStudy (anonymous):

you should get \[2xy^2+2x^2yy'-16=0\] solve for \(y'\)

OpenStudy (albert0898):

Dy/Dx = (4+16x)/(2y2x) ?

OpenStudy (anonymous):

hmm no

OpenStudy (anonymous):

\[2xy^2+2x^2yy'-16=0\\ xy^2+x^2yy'-4=0\\ xy^2+x^2yy'=4\\ x^2yy'=4-xy^2\]

OpenStudy (anonymous):

so \[y'=\frac{4-xy^2}{x^2y}\]

OpenStudy (albert0898):

How would I then solve for the 2nd derivative? Thank you

OpenStudy (anonymous):

take the derivative of that !

OpenStudy (anonymous):

requires the quotient rule, and is kind of messy, but totally doable

OpenStudy (albert0898):

Thank you :). I guess my teacher enjoys the suffering of others